OpenStudy (mtwizzle420):

Which of the following are possible side lengths for a right triangle? I will give a medal :) A.) 3,6,9 B.) 4,8,10 C.) 9,12,15 D.) 7,11,15

OpenStudy (anonymous):

Use the Pythagorean Theorem and see if they are true. \(a^2 + b^2 = c^2\) Plug in the first: \(3^2 + 6^2 = 9^2\)

OpenStudy (mtwizzle420):

Its C :) Thank you!